OpenStudy (anonymous):

Which statement most accurately explains why x^6/x^7 is equal to x^3? A. two of the five x's on top cancel with the three x's on bottom, thus x^3 B. three of the five x's on top cancel with the three x's on bottom, thus x^3 C. two of the five x's on top cancel with the two x's on bottom, leaving three x's on top, thus x^3 D. three of the five x's on top cancel with the three x's on bottom, leaving three x's on top, thus x^3 help?

OpenStudy (mertsj):

Still isn't fixed.

OpenStudy (anonymous):

ok idc I don't see the mistake no need to comment

OpenStudy (mertsj):

You have posted the problem incorrectly. In your opening statement you say that x^6/x^7=x^3 and that simply is not true.

OpenStudy (anonymous):

if I could take a picture that's exactly what it says

OpenStudy (mertsj):

Well then you need to tell your teacher that the problem is written wrong and none of the answer choices are true because the original premise is false.

OpenStudy (anonymous):

its a summer school question on the computer

OpenStudy (mertsj):

Doesn't matter. Either they have posted it wrong or you have posted it wrong. The result is the same. The original premise is false.

OpenStudy (anonymous):

re-read it because if you cant get ill never get it because your obviously way smarter den me :(

OpenStudy (mertsj):

This is what the first sentence says: \[\frac{x^6}{x^7}=x^3\]

OpenStudy (mertsj):

Is that right? That's the first sentence of the problem.

OpenStudy (anonymous):

x^6/x^7 is equal to x^3

OpenStudy (mertsj):

What does x^6 mean?

OpenStudy (anonymous):

x^6 over x^7 is equal to x^3

OpenStudy (mertsj):

x^6 means (x)(x)(x)(x)(x)(x)

OpenStudy (mertsj):

What does x^7 mean?

OpenStudy (anonymous):

listen give me a suggestion and ill go wht with you think it is. its ok if its wrong.

OpenStudy (mertsj):

What does x^7 mean?

OpenStudy (anonymous):

(x)(x)(x)(x)(x)(x)(x)

OpenStudy (mertsj):

Now write a fraction with (x)(x)(x)(x)(x)(x) on the top and

OpenStudy (mertsj):

(x)(x)(x)(x)(x)(x)(x) on the bottom

OpenStudy (mertsj):

The is x^6/x^7

OpenStudy (mertsj):



OpenStudy (mertsj):

Does yours look like that?

OpenStudy (anonymous):

no

OpenStudy (mertsj):

Then do it again and make it look like that.

OpenStudy (anonymous):

\[x ^{6}\] / \[x^{7}\]

OpenStudy (anonymous):

is equal to \[x ^{3}\]

OpenStudy (mertsj):

Do you have trouble with math?

OpenStudy (anonymous):

that's why im in summer school bro

OpenStudy (mertsj):

Do you want to know one of the reasons you have trouble with math?

OpenStudy (anonymous):

not really. but hit me with it

OpenStudy (mertsj):

Because you refuse to do what you are told.

OpenStudy (anonymous):

its exponents not math in general

OpenStudy (mertsj):

I told you to write a fraction with (x)(x)(x)(x)(x)(x) on the top and (x)(x)(x)(x)(x)(x)(x) on the bottom.

OpenStudy (mertsj):

Then I showed you how it should look

OpenStudy (mertsj):

Then I asked you if yours looked like that and you threw some totally useless gobblety gook at me that didn't help you understand the problem one bit.

OpenStudy (mertsj):

If you had just done what I told you, we would be way further on the way to solving this problem.

OpenStudy (espex):

If you're using windows 7, you can use the snipping tool to capture an image of your problem.

OpenStudy (espex):

Otherwise you can use print screen and paste it into paint. Crop it down to what you need and post.